An amplifier is set to a gain of 7.5 dB. If 7.00×10-6 W of electrical power is applied to the input, how much power will be delivered at the output?
Answer #1

here,

power input , Pi = 7 * 10^-6 W

let the power output be Po

gain = 7.5 dB = 10*log(7.5) = 10^0.75 = Po/Pi

Po = 5.62 * 7 * 10^-6

Po = 3.94 * 10^-5 W

the output power delivered is 3.94 * 10^-5 W